OpenStudy (anonymous):

Can you help me this. The ratio of adults to children at party is 4 to 5. The total number of people at the party is between 20 and 30. Write an equipvalent ration to find how many adults and children are at the party? Thnx

OpenStudy (anonymous):

the integral multiple of (4+5) =9 and between 20 and 30 is 27 hence each ratio is 27/9=3 thus number o adults =4*3=12 and children 5*3=15

OpenStudy (anonymous):

Why you pick #27?